OpenStudy (anonymous):

[6.03] Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

I would use the elimination method because when the two equations are added, the y's automatically cancel each other out. 2x + 5y = -3 -15x - 5y = -10 ---------------add -13x = -13 x = 1 now sub 1 in for x in either of the original equations to find y 2x + 5y = -3 2(1) + 5y = -3 2 + 5y = -3 5y = -3-2 5y = -5 y = -1 check... -15x - 5y = -10 -15(1) - 5(-1) = -10 -15 + 5 = -10 -10 = -10 (correct) any questions ?